I need help understanding the following:
If (-1,-4) is a point on the graph of y=f(x), what point do you know is on the graph of y=f(1/2x)?
Would the answer be (-1/2,-2) since it is 1/2 of (-1,-4)?
: I need help understanding the following:
If (-1,-4) is a point on the graph of y=f(x), what point do you know is on the graph of y=f(1/2x)?
Would the answer be (-1/2,-2) since it is 1/2 of (-1,-4)?

Answer by jojo14344(1022)

We'll see shortly:
At points (-1,-4), and consider other points (0,0):
Slope=(y[2]-y[1])/(x[2]-x[1])=(0-(-4))/(0-(-1))=4/1
Slope=4, "m"
Thru points (-1,-4):
y=mx+b, slope intercept form
-4=(-1)4+b
-4+4=0=b, y-intercept=0:
Then, the line eqn follows: y=4x
drawing(300,300,-5,5,-5,5,grid(1),graph(300,300,-5,5,-5,5,4x),circle(-1,-4,.20))-->(y)=4x ---> points (-1,-4)
If (y)=(1/2)(x), then,
drawing(300,300,-5,5,-5,5,grid(1),graph(300,300,-5,5,-5,5,(1/2)x),circle(-1,-1/2,.20)) --->(y)=(1/2)x ----> points (-1,-1/2), ANSWER
Thank you,
Jojo